Two point targets are separated by 3m in slant range, (a) What radar bandwidth is required to resolve the two targets? (b) If the radar is a monotone pulse radar, what should be the maximum transmitted pulse duration? 


1
Expert's answer
2021-01-27T04:55:18-0500


Distance between the targets = 3 m

a)

The radar Range Resolution (d) is given by;

"d=\frac{c}{2B}"

"\frac{3 \times10^8}{ 3\times B}= 3"

"B = 5 \times10^7 Hz = 50MHz"

b) The pulse duration of radar pulse is given by:

"\tau =\frac{1}{B}"

"\frac{1}{(5\times 10^7)} = 0.2\times 10^7 s"